OpenStudy (anonymous):

Please help fast!! Solve the equation by first using a Sum-to-Product formula. (Enter your answers as a comma-separated list. Let k be any integer. Round terms to three decimal places where appropriate.) cos 5θ − cos 7θ = 0

OpenStudy (destiny23):

cos(a - b) - cos(a + b) => cos(a)cos(b) + sin(a)sin(b) - (cos(a)cos(b) - sin(a)sin(b)) => cos(a)cos(b) - cos(a)cos(b) + sin(a)sin(b) + sin(a)sin(b) => 2sin(a)sin(b) cos(5t) - cos(7t) => cos(6t - t) - cos(6t + t) => 2 * sin(6t) * sin(t) 2 * sin(6t) * sin(t) = 0 sin(6t) = 0 6t = pi * k t = (pi/6) * k sin(t) = 0 t = pi * k Most general answer: t = (pi/6) * k